OpenStudy (anonymous):

Can someone help me with this Integral? integral from 0 to 5 (x+1/(sqrt(x+4)) The correct answer is 20/3. P.S. I tried u- substitution but it didn't work for me.

hartnn (hartnn):

is it \(\large \int \limits_0^5 \dfrac{x+1}{\sqrt{x+4}}dx\) the substi. of \(u=\sqrt{x+4}\) should work.

OpenStudy (anonymous):

ok I'll try it again. I used u=x+4

hartnn (hartnn):

ohh...try \(u=\sqrt {x+4}\) not quite difficult :)

hartnn (hartnn):

so, did you get it ? or still trying ?

OpenStudy (anonymous):

still trying to work it out.. trying to find x

hartnn (hartnn):

find x ? in terms of u ? see whether u get x= u^2 -4

OpenStudy (anonymous):

then is du= 1/2sqrt(x+4) ?

hartnn (hartnn):

yes.

OpenStudy (anonymous):

ok i got it thanks!

hartnn (hartnn):

happy to help :) welcome ^_^